OpenStudy (anonymous):

6^5a^8b^3/(6ab^2c)^4=

OpenStudy (anonymous):

try to eliminate the comon terms

OpenStudy (mysesshou):

\[\frac{6^5a^8b^3}{(6ab^2c)^4}\] \[\frac{6^5a^8b^3}{(6^4a^4b^8c^4)}\]

OpenStudy (mysesshou):

\[\frac{6^5a^8b^3}{6^4a^4b^8c^4}\] 6^5/6^4 = 6 a^8/a^4 = a^4 b^3/b^8 = 1/b^5 \[\frac{6a^4}{b^5c^4}\] i think.